OpenStudy (anonymous):

solve:x^5-1 How to solve this equation?

OpenStudy (anonymous):

i would have helped u. if you weren't so Mouthly protesting in Group chat

OpenStudy (anonymous):

take x^5-1=0, x^5=1.. now by the complex number you can solve it...take x^5=1+0i x^5=sin(pi/2)+icos(pi/2) then take 5th power and use De morves theorem ...there will be 5 sollution in complex numbers